# The Generalization Ridge: Information Flow in Natural Language Generation

### v1

@@ Abstract
Transformer-based language models have achieved state-of-the-art performance in natural language generation (NLG) tasks, yet their internal mechanisms for synthesizing task-relevant information remain insufficiently understood. While prior studies suggest that intermediate layers often yield more generalizable representations than final layers, how this generalization ability emerges and propagates across layers during training remains unclear. To address this gap, we propose InfoRidge, an information-theoretic framework, to characterize how predictive information—the mutual information between hidden representations and target outputs—varies across depth. Estimating this quantity enables us to trace the flow of task-relevant information throughout the model during training. Our experiments across various models and datasets reveal a consistent non-monotonic trend: predictive information peaks in upper-middle layers—forming a generalization ridge—before declining in final layers, reflecting a transition between generalization and memorization. To further investigate this phenomenon, we introduce residual scaling coefficients—trainable scalar parameters applied to each residual block—which serve as functional probes for assessing the relative importance of individual transformer layers. These coefficients reveal that, under distribution shift, models downweight final layers and increasingly rely on ridge layers, highlighting their role in generalization. Together, these findings offer new insights into the internal mechanisms of transformers and underscore the critical role of intermediate layers in supporting generalization.
@@ 1 Introduction
Transformer-based language models have achieved remarkable performance in natural language generation (NLG) tasks such as summarization, machine translation, and dialogue generation vaswani2017attention ; dong2022survey . Nevertheless, we lack a rigorous understanding of how these models acquire and synthesize task-relevant information during training.
@@
A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance liu2019linguistic ; voita2019bottom ; ansuini2019intrinsic ; ahrens2023read ; uselis2025intermediate . In language models, intermediate layers often encode richer semantic and more robust features than final layers fan2024not ; jin2024exploring ; skean2025layer . However, questions still remain: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
@@
To investigate these questions, we propose InfoRidge, an information-theoretic framework, to analyze information flow in language models. Building on matrix-based mutual information estimation giraldo2014measures , our approach quantifies how predictive signals transform across layers during training. We center our analysis on two complementary quantities. The first is the predictive information at each layer, defined as the mutual information $I(Z_{\ell};Y)$ between the hidden representation $Z_{\ell}$ at layer $\ell$ and the next-token label $Y$, which reflects how much task-relevant information is preserved. The second is the incremental information gain, denoted as $I(\Delta Z_{\ell};Y)$, where $\Delta Z_{\ell}$ is the residual changes between successive layer $\ell$ and $\ell-1$. This measures the additional predictive information introduced by each transformer block.
@@
Using these tools, we uncover a non-monotonic trend: predictive information rises through early and middle layers, peaks in the upper-middle, and then declines in later layers. We term this phenomenon the generalization ridge, indicating a structural division of labor across depth. Intermediate layers appear to specialize in encoding abstract, generalizable features, while later layers increasingly reorient representations toward task-specific alignment and memorization. By further analyzing incremental information gain, we find that certain middle layers introduce the largest increases in predictive information, marking them as key contributors to the emergence of the ridge. Such layer-wise specialization offers insight into how generalization and memorization are distributed within the network.
@@
To further validate this interpretation, we introduce residual scaling coefficients—learnable scalar parameters $\beta_{\ell}$ applied to each residual block—while keeping all other model weights fixed, drawing inspiration from prior work on layer-wise adaptation liu2019self ; menghani2024laurel . A higher $\beta_{\ell}$ value indicates greater reliance on the corresponding layer’s output during prediction. These coefficients act as functional probes, revealing how the model redistributes layer-wise importance under different data distribution. Under in-distribution training, deeper layers retain higher residual weights, reflecting the model’s reliance on memorized, task-specific features. When evaluated under distribution shift, models reduce reliance on late layers and increase reliance on the ridge, further supporting its role in generalization.
@@
To understand the formation of the generalization ridge, we analyze both attention patterns and model capacity. Attention flow analysis shows that ridge layers attend more directly to semantically informative tokens, suggesting that the ridge is not only a statistical peak but also functionally aligned with meaningful computation. In contrast, deeper layers increasingly focus on positional tokens, consistent with a shift toward memorization. Beyond functional behavior, we also investigate the conditions under which the ridge emerges. Our depth ablation results show that the ridge only emerges beyond a certain depth threshold. Below this threshold, predictive information increases monotonically, indicating that sufficient capacity is a prerequisite for generalization ridge to emerge.
@@
  1. 1.
    We propose InfoRidge, an information-theoretic framework, to quantify both predictive information and incremental information gain across layers, offering a dynamic view of how task-relevant signals emerge and evolve during training.
  2. 2.
    Through this framework, we identify the generalization ridge—a mid-layer peak in predictive information—revealing a transition between generalization and memorization.
  3. 3.
    We propose residual scaling coefficients as functional probes for layer importance, uncovering a shift in reliance from final layers to intermediate layers under distribution shift, consistent with the ridge’s role in supporting generalization.
These findings suggest that transformer depth reflects functional stratification: intermediate layers encode core abstractions for generalization, while final layers increasingly tailor representations for memorization—often at the cost of robustness. This insight informs the design of more interpretable and adaptable architectures under distribution shift.
@@ 2 Related Work
Understanding how information is encoded and transformed across layers in deep models has been widely studied through probing classifiers alain2016understanding , attention flow analysis vig2019analyzing , and information-theoretic tools such as the information bottleneck shwartz2017opening , mutual information estimation goldfeld2019estimating , and matrix-based entropy methods giraldo2014measures . These methods provide various lenses to quantify representational capacity, abstraction, and invariance across layers.
A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance ansuini2019intrinsic ; yosinski2014transferable ; uselis2025intermediate ; ahrens2023read ; ando2023use . In language models, mid-depth layers tend to capture richer semantic or robust features than output layers liu2019linguistic ; voita2019bottom ; jin2024exploring ; fan2024not . These findings challenge the assumption that deeper layers always yield better representations, revealing instead a peak in representational utility at intermediate depths. This pattern holds across settings such as transfer learning yosinski2014transferable , continual learning ahrens2023read , and out-of-distribution generalization uselis2025intermediate . Recent work by Skean et al. skean2025layer proposes a unified framework that evaluates representation quality using entropy, curvature, and invariance. Collectively, these results highlight the importance of intermediate layers for abstraction and generalization, motivating our analysis of layer-wise information dynamics through an information-theoretic lens. Differently, we focus on information flow for next token generation, specifically examining mutual information between the last input token and the target label. Our goal is to understand how generalization and memorization dynamics evolve during training from information-theoretic perspective.
@@ 3 InfoRidge: Information Estimation Framework
Figure 1: Overview of InfoRidge. (1) Extract internal representations at each layer and compute residual changes between successive layers. (2) Extract the target token embedding. (3) Estimate predictive information and incremental information gain.
@@
Prior work has shown that internal representations in deep neural networks tend to align most closely with the true data distribution at an intermediate layer he2024information . By employing the Wasserstein distance villani2008optimal , this alignment is shown to reach a minimum at a specific depth—referred to as the generalization funnel layer. At this point, the Min Wasserstein Generalization Bound he2024information ensures that the upper bound on the generalization gap—defined as the expected difference between the population and empirical risks—is minimized. This highlights the critical role of intermediate layers in supporting generalization.
@@
Building on this insight, we hypothesize that there exists a specific intermediate transformer layer that encodes the most generalizable representations for next-token prediction, characterized by maximal mutual information with the target label.
@@
We apply the matrix-based framework giraldo2014measures to estimate mutual information. Let $\mathcal{U}$ be a random variable, from which we draw a set of vectors $U=\{\mathbf{u}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$. A positive-definite Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is computed using a Gaussian kernel $\kappa$ with bandwidth set to 1 and the matrix is then trace-normalized to satisfy $\operatorname{tr}(G_{U})=1$. The matrix-based Rényi entropy (with order $\alpha=1$) is then given by:
@@ 4 Experimental Setup
We evaluate four models: GPT–2 Small (117M) radford2019language , GPT–2 Medium (345M) radford2019language , Qwen–2.5 0.5B qwen2 , and LLaMA 3.1 8B meta_llama_3.1_8b . All models are fine-tuned on NLG tasks. Each model shares weights between the input token embedding and the output language modeling head.
@@
We assess model behavior across three tasks reformulated as NLG problems: CLUTRR (sinha2019clutrr, ), a relational reasoning benchmark; ECQA (aggarwal2021explanations, ), a commonsense QA benchmark; and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise. Dataset statistics and implementation details are provided in Appendix B and C.
[…]
@@ 5.1 Predictive Information: Information Peaks at Intermediate Layers
We investigate how predictive information—defined as the mutual information between hidden representations and target labels—evolves across transformer layers. Specifically, we calculate the matrix–based mutual information $I(Z_{\ell};Y)$ at each transformer layer $\ell$, capturing the amount of label-relevant information encoded during forward propagation.
@@
Figure 2 tracks the trajectory of the predictive information $I(Z_{\ell};Y)$ between the hidden representation at depth $\ell$ and the target label $Y$ throughout training, while Table 1 reports the downstream accuracy obtained when we early exit after a given layer. [^1] The in distribution split is generated with $K_{\text{id}}{=}13$; out of distribution splits use a uniformly–sampled $K_{\text{ood}}\in\{5,\dots,25\}\setminus\{13\}$. Additional results are presented in Appendix E.
@@
Table 1: Information Dynamics and Layer-wise Performance (GPT-2 Small, Synthetic). OOD performance declines beyond the generalization ridge.
Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist. Out-Dist. Layer 1 0.0508 0.00 0.00 0.00 Layer 2 0.0513 0.00 0.00 0.00 Layer 3 0.0619 0.00 0.00 0.00 Layer 4 0.0822 0.00 0.00 0.00 Layer 5 0.0894 0.00 0.00 0.00 Layer 6 0.0989 0.00 0.00 0.00 Layer 7 0.1292 0.70 0.00 2.00 Layer 8 0.2431 18.35 0.00 38.90 Layer 9 0.2810 19.15 1.00 40.30 Layer 10 0.3209 62.45 71.90 53.50 Layer 11 0.3209 72.65 99.90 40.40 Layer 12 0.1402 71.45 100.00 38.60
Information Peak (intermediate layers). $I(Z_{\ell};Y)$ continues to rise through the mid-to-upper layer, typically peaking before the final few blocks. For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.32$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $72\%$ ID accuracy and $53\%$ OOD accuracy. These layers appear to play a critical role in synthesizing abstract features that are essential for generalization.
@@
Discussion on Model Capacity Threshold
@@
(c) Full model:Generalization Ridge
@@
We empirically vary the depth of GPT-2 Small models fine-tuned on the CLUTRR dataset to examine how model capacity shapes information dynamics, shown in Figure 3. When the model is truncated to 8 layers, $I(Z;Y)$ increases monotonically—indicating insufficient capacity to develop an information peak. Upon increasing the depth to 9 layers, a shallow peak emerges, signifying the threshold at which the model begins to distinguish generalized features from memorized signals. The full 12-layer model exhibits a clear peak followed by a decline, confirming that generalization ridge emerges only beyond a certain capacity threshold. These results underscore the role of architectural depth in information dynamics and sufficient capacity is required for generalization ridge to emerge.
Discussion on Overfitting Scenario
@@
To probe overfitting dynamics, we intentionally fine-tuned the model beyond the optimal point. In Figure 4, we observe that $I(Z_{\ell};Y)$ rises again in the final layers—departing from the typical compression phase. This reversal may suggest that the model begins to memorize shortcuts or redundant label-specific noise rather than learning useful information.

### v2

@@ Abstract
Transformer-based language models have achieved state-of-the-art performance in natural language generation (NLG), yet their internal mechanisms for synthesizing task-relevant information remain insufficiently understood. While prior studies suggest that intermediate layers often yield more generalizable representations than final layers, how this generalization ability emerges and propagates across layers during training remains unclear. To address this gap, we propose InfoRidge, an information-theoretic framework, to characterize how predictive information—the mutual information between hidden representations and target outputs—varies across depth. Estimating this quantity enables us to trace the flow of task-relevant information throughout the model during training. Our experiments across various models and datasets reveal a consistent non-monotonic trend: predictive information peaks in upper-middle layers—forming a generalization ridge—before declining in final layers, reflecting a transition between generalization and memorization. To further investigate this phenomenon, we introduce residual scaling coefficients—trainable scalar parameters applied to each residual block—which serve as functional probes for assessing the relative importance of individual transformer layers. These coefficients reveal that, under distribution shift, models downweight final layers and increasingly rely on intermediate layers, highlighting their role in generalization. Together, these findings offer new insights into the internal mechanisms of transformers and underscore the critical role of intermediate layers in supporting generalization.
@@ 1 Introduction
Transformer-based language models generate text by predicting tokens autoregressively, and they have achieved remarkable performance across a wide range of natural-language uses (Vaswani et al., 2017; Dong et al., 2022). Nevertheless, we lack a rigorous understanding of how these models acquire and synthesize task-relevant information during training.
@@
A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance (Liu et al., 2019b; Voita et al., 2019; Ansuini et al., 2019; Ahrens et al., 2023; Uselis & Oh, 2025). In language models, intermediate layers often encode richer semantic and more robust features than final layers (Fan et al., 2024; Jin et al., 2024; Skean et al., 2025). However, questions still remain: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
@@
To investigate these questions, we propose InfoRidge, an information-theoretic framework, to analyze information flow in language models. Building on matrix-based mutual information estimation (Giraldo et al., 2014), our approach quantifies how predictive signals transform across layers during training. We center our analysis on two complementary quantities: predictive information, defined as the mutual information $I(Z_{\ell};Y)$ between the hidden representation $Z_{\ell}$ at layer $\ell$ and the next-token label $Y$, reflecting how much task-relevant information is preserved; incremental information gain, denoted as $I(\Delta Z_{\ell};Y)$, where $\Delta Z_{\ell}$ is the residual changes between successive layer $\ell$ and $\ell-1$, measuring the additional predictive information introduced by each transformer block.
@@
Using InfoRidge, we uncover a non-monotonic trend: predictive information rises through the early and middle layers, peaks in the upper-middle layers, and then declines in the later layers. We name this phenomenon the generalization ridge, where the model encodes the more generalizable task-relevant information. This ridge marks a structural division of labor: intermediate layers concentrate generalizable features that transfer across distributions, while later layers increasingly specialize in task-specific memorization. Incremental information gain further shows that ridge layers introduce the largest increases in predictive information, marking them as key contributors to the emergence of the ridge. This analysis directly connects the information peak to generalizable behavior and clarifies how generalization and memorization are distributed across depth.
@@
To further validate this interpretation, we introduce residual scaling coefficients—learnable scalar parameters $\beta_{\ell}$ applied to each residual block—while keeping all other model weights frozen, drawing inspiration from prior work on layer-wise adaptation (Liu et al., 2019a; Menghani et al., 2024). A higher $\beta_{\ell}$ value indicates greater reliance on the corresponding layer’s output during prediction. These coefficients act as functional probes, revealing how the model redistributes layer-wise importance under different data distribution. Under in-distribution training, deeper layers retain higher residual weights, reflecting the model’s reliance on memorized, task-specific features. When evaluated under distribution shift, models have lower reliance on the final layer and shifts importance toward the intermediate layers.
@@
To understand the formation of the generalization ridge, we analyze both attention patterns and model capacity. Attention analysis shows ridge layers attend to tokens that capture broadly useful features, aligning with the information peak. Beyond attention patterns, we also investigate the conditions under which the ridge emerges. Our depth ablation results show that the ridge only emerges beyond a certain depth threshold. Below this threshold, predictive information increases monotonically, indicating that sufficient capacity is a prerequisite for generalization ridge to emerge.
@@
1. Our work tracks the evolution of predictive information throughout training, establishing a clear connection between predictive information flow and generalization. It reveals a non-monotonic peak in the middle layers, which we refer to as the generalization ridge. This pattern reflects a meaningful transition in representational focus and aligns with stronger generalization behavior.
2. We introduce InfoRidge, an information-theoretic framework that applies matrix-based mutual information estimation to autoregressive language models to analyze information flow.
3. We introduce residual scaling coefficients as trainable indicators of how models shift representational focus across layers during training, providing a causal, adaptive measure of model behavior under distribution shifts.
@@ 2 Related Work
Understanding how information is encoded and transformed across layers has been studied through probing classifiers (Alain & Bengio, 2016), attention flow (Vig & Belinkov, 2019), and information-theoretic approaches such as the information bottleneck (Shwartz-Ziv & Tishby, 2017), mutual information estimation (Goldfeld, 2019), and matrix-based entropy (Giraldo et al., 2014), offering different lenses to quantify representational capacity, abstraction, and invariance across layers.
A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance (Ansuini et al., 2019; Yosinski et al., 2014; Uselis & Oh, 2025; Ahrens et al., 2023; Ando et al., 2023). In language models, mid-depth layers tend to capture richer semantic or robust features than output layers (Liu et al., 2019b; Voita et al., 2019; Jin et al., 2024; Fan et al., 2024). Transformer representations have long been observed to follow a structured progression from syntactic to semantic information, as shown by classical probing studies on linguistic knowledge and the reconstruction of the NLP pipeline (Liu et al., 2019b; Tenney et al., 2019). These findings challenge the assumption that deeper layers always yield better representations.
This pattern holds across settings such as transfer learning (Yosinski et al., 2014), continual learning (Ahrens et al., 2023), and out-of-distribution generalization (Uselis & Oh, 2025). Furthermore, recent work has evaluated representation quality using entropy, curvature, and invariance (Skean et al., 2025), while other studies have examined embedding drift and representational geometry (Merchant et al., 2020; Dar et al., 2022), analyzed memorization and factual recall (Haviv et al., 2022; Yu et al., 2023), and introduced causal perspectives on layer importance through mediation analysis and targeted interventions (Vig et al., 2020; Meng et al., 2022). Training-dynamics studies investigated how earlier models develop and refine semantic features across depth, providing an additional perspective on the evolution of layer-wise representations (Merchant et al., 2020; Kumar et al., 2023). Correlational probes (e.g., linear probes (Alain & Bengio, 2016)) measure only whether a feature can be decoded from a representation, which reflects correlation but not causal influence. In contrast, causal methods intervene on internal activations to test how changing a component alters the model’s prediction, thereby identifying true causal contribution rather than mere feature presence. Our residual-scaling approach aligns with this causal perspective at the layer level.
However, the underlying causes and functional role of this phenomenon remain only partially understood, motivating further investigation. Our work addresses this gap by tracing the evolution of predictive information throughout training and establishing a clear connection between predictive information flow and generalization. We reveal a consistent non-monotonic peak in the middle layers—termed the generalization ridge—that reflects a meaningful transition in representational focus and aligns with stronger generalization behavior. Additionally, unlike prior work focused on classification tasks, we extend the analysis of information flow to generation tasks by quantifying the mutual information between hidden states and the next token. This enables us to understand how generalization and memorization dynamics evolve for next-token generation setting during training from an information-theoretic perspective.
@@ 3 InfoRidge: Information Estimation Framework
Figure 1: Overview of InfoRidge. (1) Step 1: Extract internal representations at each layer and compute residual changes between successive layers. (2) Step 2: Extract the target token embedding. (3) Step 3: Using these representations, we estimate the layer-wise predictive information $I(Z_{\ell};Y)$ and the incremental information gain $I(\Delta Z_{\ell};Y)$, which jointly characterize how information evolves across depth.
@@
Prior work has shown that internal representations in deep neural networks tend to align most closely with the true data distribution at an intermediate layer (He et al., 2024). By employing the Wasserstein distance (Villani et al., 2008), this alignment is shown to reach a minimum at a specific depth—referred to as the generalization funnel layer. At this point, the Min Wasserstein Generalization Bound (He et al., 2024) ensures that the upper bound on the generalization gap—defined as the expected difference between the population and empirical risks—is minimized. This highlights the critical role of intermediate layers in supporting generalization.
Research Question.
Despite the insight from prior studies of deep neural networks in classification, it remains unclear whether this generalizes to Transformer-based language models in natural language generation (NLG). This motivates our central question: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
@@
Building on the insight, we hypothesize that there exists a specific intermediate transformer layer that encodes the most generalizable representations for next-token prediction, characterized by maximal mutual information with the target label.
@@
We apply the matrix-based framework (Giraldo et al., 2014) to estimate mutual information. Let $\mathcal{U}$ be a random variable, from which we draw a set of vectors $U=\{\mathbf{u}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$. A positive-definite Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is computed using a Gaussian kernel $\kappa$ with bandwidth set to 1 and the matrix is then trace-normalized to satisfy $\operatorname{tr}(G_{U})=1$. The matrix-based Rényi entropy (with order $\alpha=1$) is then given by:
@@ 4 Experimental Setup
We evaluate four models: GPT–2 Small (117M) (Radford et al., 2019), GPT–2 Medium (345M) (Radford et al., 2019), Qwen–2.5 0.5B (Yang et al., 2024), and LLaMA 3.1 8B (Meta AI, 2024). All models are fine-tuned on NLG tasks. Each model shares weights between the input token embedding and the output language modeling head.
@@
We assess model behavior across three tasks casted into NLG problems: CLUTRR (Sinha et al., 2019), a relational reasoning benchmark; ECQA (Aggarwal et al., 2021), a commonsense QA benchmark; and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise. Dataset and implementation details are in Appendix B and C.
Synthetic Arithmetic Dataset Construction. We construct a synthetic dataset to separate signal learning from noise memorization. Each input is a sequence of 10 elements, where the signal follows an arithmetic progression modulo $K$, computed as $s_{t}=(s_{0}+t\cdot d)\bmod K$ with $s_{0}\in[0,K{-}1]$ and $d\in[1,K{-}1]$. Each element takes the form S{signal}_N{noise}, where noise is sampled from $\mathcal{U}_{\text{int}}(0,\texttt{noise\_range-1})$ ($\mathcal{U}_{\text{int}}$ denotes the uniform distribution). The model is trained to predict the signal value of the final (10th) element using the preceding elements as input context. For example, with $K=5$, $s_{0}=1$, and $d=2$, a sample input might be S1_N42 S3_N77...S2_N37, with the target signal being 4. By varying $K$, we can induce structured distribution shifts.
@@ 5.1 Predictive Information: Information Peaks at Intermediate Layers
We investigate how predictive information—defined as the mutual information between hidden representations and target labels—evolves across the depth of transformer models. Specifically, for each layer $\ell$, we compute the matrix-based mutual information $I(Z_{\ell};Y)$ between the hidden state $Z_{\ell}$ and the next-token ground truth $Y$. This quantity measures how much task-relevant signal is retained in the representation as it propagates through the network. By tracing $I(Z_{\ell};Y)$ across layers, we obtain a layer-wise trajectory of information flow, which reveals not only where predictive content is preserved but also how it is transformed or diminished as the model processes.
@@
Figure 2 tracks the trajectory of the predictive information $I(Z_{\ell};Y)$ between the hidden representation at depth $\ell$ and the target label $Y$ throughout training, while Table 1 reports the downstream accuracy obtained when we early exit after a given layer. [^1] The in distribution split is generated with $K_{\text{id}}{=}13$; out of distribution splits use a uniformly–sampled $K_{\text{ood}}\in\{5,\dots,25\}\setminus\{13\}$. Additional results are in Appendix E.
@@
Table 1: Information Dynamics and Layer-wise Performance (GPT-2-S, Synthetic). OOD performance declines beyond the generalization ridge.
Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist. Out-Dist. Layer 1 0.0782 0.00 0.00 0.00 Layer 2 0.0776 0.00 0.00 0.00 Layer 3 0.0907 0.00 0.00 0.00 Layer 4 0.1156 0.00 0.00 0.00 Layer 5 0.1222 0.00 0.00 0.00 Layer 6 0.1275 0.00 0.00 0.00 Layer 7 0.1552 0.75 0.00 1.50 Layer 8 0.2594 19.45 0.00 40.70 Layer 9 0.2937 20.60 0.60 42.00 Layer 10 0.3295 64.25 72.90 56.90 Layer 11 0.3296 71.65 100.00 42.90 Layer 12 0.1501 70.45 100.00 41.40
@@
Information Peak (intermediate layers). $I(Z_{\ell};Y)$ continues to rise through the mid-to-upper layer, typically peaking before the final few blocks. For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.33$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $73\%$ ID accuracy and $57\%$ OOD accuracy. These layers appear to play a critical role in synthesizing abstract features that are essential for generalization.
@@
Generalization Ridge Emerges Beyond a Depth Threshold
@@
(c) Full model: Gen Ridge
@@
We empirically vary the depth of GPT-2 Small models fine-tuned on the CLUTRR dataset to examine how model capacity shapes information dynamics, shown in Figure 3. When the model is truncated to 8 layers, $I(Z;Y)$ increases monotonically—indicating insufficient capacity to develop an information peak. Upon increasing the depth to 9 layers, a shallow peak emerges, signifying the threshold at which the model begins to distinguish generalized features from memorized signals. The full 12-layer model exhibits a clear peak followed by a decline, confirming that the generalization ridge emerges only beyond a certain capacity threshold. These results underscore the role of architectural depth in information dynamics and sufficient capacity is required for the generalization ridge to emerge.
Overfitting Adds Memorization in Final Layers
@@
To probe overfitting dynamics, we intentionally fine-tuned the model beyond the optimal point. In Figure 4, we observe that the Predictive Information $I(Z_{\ell};Y)$ rises again in the final layers—departing from the typical compression phase expected at this stage. Before overfitting, the top layers largely behave as pretrained decoders, focusing on surface-level patterns such as token co-occurrence, syntactic templates, or corpus biases, which are not strongly aligned with the task label—hence their initially low $I(Z_{\ell};Y)$. After overfitting, however, the later-layer rise reflects a shift toward memorization, where the model begins to encode superficial shortcuts or redundant label-specific noise rather than useful task information.
Semantically Important Attention Peaks Where Information Peaks
This discussion aims to provide an intuitive illustration of our generalization ridge hypothesis, highlighting how attention to task-relevant signal tokens shifts across layers (the generalizable information). Specifically, we computed average signal attention across layers, identified the layer with peak signal attention, and compared it to final-layer signal and last-token attention, results are in Table 2. Signal tokens are defined as task-relevant tokens that the model must focus on to solve the task—for the synthetic dataset, these are tokens that appear after the character ‘S’; for CLUTRR, kinship-related entities; and for ECQA, the token corresponding to the correct answer option. This allows us to quantify where in the network attention to semantically important information is concentrated. Our findings show a pattern that supports our hypothesis: (1) Mid-to-late layers peak in signal attention, coinciding with the predictive information ridge (Figure 2), indicating where generalizable representations are strongest. (2) Final layers show reduced attention to signal tokens, suggesting a shift toward memorization to specific data point rather than predictive information abstraction. For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2458) but drops to 0.0066 in the final layer, where last-token attention dominates (0.1697). Additional results are in Appendix F.
Table 2: Average attention statistics: (1) average attention scores over all tokens, (2) average attention to signal tokens, (3) the maximum signal attention and its corresponding layer, (4) signal token attention in the final layer and (5) last token attention in the final layer.
Model Dataset Avg. Attn (All) Avg. Attn (Signal) Layer w/ Highest Avg. Signal Final Avg. Signal Final Avg. Last
GPT-2 Small Synthetic 0.0227 0.0483 10 (0.0758) 0.0573 0.0400
GPT-2 Medium CLUTRR 0.0086 0.0155 19 (0.0453) 0.0147 0.0889
Qwen-2.5-0.5B ECQA 0.0225 0.0257 17 (0.2458) 0.0066 0.1697
LLaMA-3.1-8B ECQA 0.0220 0.0379 21 (0.1276) 0.0168 0.3904
*We remove the first token attention score to mitigate attention sink effects.